How do I solve this system of equations?
y=2x+3
y=-2x+7

Answer by blwinbbbles(106)

You would use substitution
if y = 2x + 3 and y = -2x + 7 then you can say:
2x + 3 = -2x + 7 add 2x to both sides
4x + 3 = 7 subtract 3 from both sides
4x = 4 divide by 4
x = 1
so if x = 1
then go back to one of the original equations and put 1 in place of x
therefore:
y = 2(1) + 3
y = 2 + 3
y = 5
your cordinate is (1,5)